OpenStudy (anonymous):

Which equation describes the line with a slope of 2 that contains the point (4, -3)?

OpenStudy (anonymous):

y - 4 = 2(x + 3) 2(y - 3) = x + 4 y + 3 = 3(x - 4) 2(y + 4) = x + 3

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

another point slope question, did you try doing this yet?

OpenStudy (anonymous):

yes i got C but im not sure

OpenStudy (dtan5457):

y + 3 = 3(x - 4) has everything right except that the slope there is 3 and not 2, are you sure the question says the slope is 2?

OpenStudy (anonymous):

yes i am sure

OpenStudy (dtan5457):

y - 4 = 2(x + 3)

OpenStudy (dtan5457):

that takes a lot of switching and stuff to get or you can simply see that x =2x aka the slope is 2 here

OpenStudy (anonymous):

yes i notice that now